Verify that the function u(x,y,z)=1/(√x^2+y^2+z^2) is a solution of the three-dimensional Laplace equation Uxx + Uyy +Uzz = 0.



1
Expert's answer
2021-11-29T13:42:23-0500

"u(x,y,z)=\frac{1}{\sqrt{x^2+y^2+z^2}}"


"u_x=-\frac{x}{(x^2+y^2+z^2)^{3/2}}"

"u_{xx}=-\frac{1}{(x^2+y^2+z^2)^{3/2}}+\frac{3x^2}{(x^2+y^2+z^2)^{5/2}}=\frac{3x^2-(x^2+y^2+z^2)}{(x^2+y^2+z^2)^{5/2}}=\frac{2x^2-y^2-z^2}{(x^2+y^2+z^2)^{5/2}}"

"u_{yy}= \frac{2y^2-x^2-z^2}{(x^2+y^2+z^2)^{5/2}}"

"u_{zz}= \frac{2z^2-x^2-y^2}{(x^2+y^2+z^2)^{5/2}}"


"u_{xx}+u_{yy}+u_{zz}= \frac{2x^2-y^2-z^2}{(x^2+y^2+z^2)^{5/2}}+ \frac{2y^2-x^2-z^2}{(x^2+y^2+z^2)^{5/2}}+ \frac{2z^2-x^2-y^2}{(x^2+y^2+z^2)^{5/2}}=0"


So, "u(x,y,z)=\frac{1}{\sqrt{x^2+y^2+z^2}}" is a solution of the three-dimensional Laplace equation "u_{xx}+u_{yy}+u_{zz}=0"